# Video Generation Models Are Good Latent Reward Models

### v2

@@ 1 Introduction
Recent video generation models [lin2024open, kong2024hunyuanvideo, wan2025, gao2025seedance] have demonstrated remarkable visual fidelity, producing photorealistic content across diverse applications [liu2024video, polyak2024movie, henschel2025streamingt2v]. Nevertheless, aligning these models with complex human preferences, particularly concerning motion quality [liu2025improving], physical plausibility [bansal2025videophy], and prompt following [Cheng_2025_ICCV], remains a fundamental challenge [zeng2024dawn]. Reward feedback learning (ReFL) [xu2023imagereward, prabhudesai2023aligning, clarkdirectly2024] has emerged as a promising way for preference alignment in image generation. Nevertheless, its direct application to video generation encounters critical computational and optimization barriers that fundamentally limit its applicability.
@@
Typical ReFL approaches rely on outcome-based reward models built upon Vision-Language Models (VLMs) [unifiedreward, xu2023imagereward, he2024videoscore, liu2025improving, he2025videoscore2]. As shown in Fig. 2, such outcome-based reward models focus on videos in RGB space, which imposes three interconnected limitations due to their inherent post-hoc nature for diffusion-based video generation models (VGMs). Initially, the RGB inputs for VLMs require near-complete denoising, which introduces significant evaluation delay and drastically slows down training iterations. Furthermore, this necessity leads to a GPU memory bottleneck, as backpropagation through the VAE decoder for all video frames frequently causes GPU memory overflow. Ultimately, the issue of insufficient supervision occurs because rewards are only applied at the final steps, failing to directly guide early generation stages when structure and motion are formed. Although some recent methods [wu2024deep, Guo_2025_ICCV] attempt to distribute gradient updates via gradient stopping or trajectory shortcuts, they still depend on fully denoised frames and costly VAE decoding for reward model construction. ContentV [lin2025contentv] bypasses VAE decoding by optimizing only the first frame, yet sacrifices holistic video quality assessment, while DOLLAR [ding2025dollar] employs a VLM-based latent reward model but lacks fine-grained timestep-wise supervision.
Alternatively, VGMs themselves have shown potential as reward sources, offering process-level supervision throughout the denoising trajectory. For instance, LPO [zhang2025diffusion] pioneered using diffusion models as noise-aware latent reward models for image generation, while VideoAlign [liu2025improving] briefly mentioned using video generation models for inference-time guidance. However, the use of video generation models as train-time reward models remains largely unexplored. We posit that VGMs possess unique properties for reward modeling: (1) inherent noise-aware feature extraction at arbitrary denoising step, (2) sensitivity to generation artifacts, and (3) native support for full-sequence processing without frame sampling. Through experimental analysis, we verify that the latent features of VGMs are suitable for determining the quality of videos. Yet a critical question remains: how to repurpose VGMs into effective reward models that compress comprehensive spatiotemporal features while maintaining noise-level sensitivity.
@@ 2.1 Visual Reward Feedback Learning
Reward feedback learning has proven effective for improving visual generation quality. Early works [xu2023imagereward, clarkdirectly2024, prabhudesai2023aligning] finetune diffusion models using differentiable reward signals from human preferences, but face a depth-efficiency dilemma: step-by-step backpropagation through extensive denoising steps incurs prohibitive memory costs, while training only final steps fails to optimize low-level objectives like symmetry. Recent approaches [wu2024deep, Guo_2025_ICCV] mitigate this through gradient stopping or trajectory-preserving shortcuts. However, these image-focused methods encounter severe computational barriers in video generation due to multi-frame synthesis and gradient-enabled VAE decoding overhead. ContentV [lin2025contentv] optimizes only the first frame but cannot capture comprehensive video quality. Critically, almost existing approaches require pixel-space rewards, necessitating VAE decoding. While DOLLAR [ding2025dollar] leverages a VLM as a latent reward model for distillation tasks to avoid costly VAE decoding, it lacks timestep-wise optimization capability. In this paper, we propose a timestep-wise latent reward model operating directly in latent space, enabling gradient backpropagation at arbitrary timesteps while maintaining computational tractability.
@@ 2.2 Reward Models for Video Generation
Reward models for RLHF [ouyang2022training] are typically trained on human preferences [sun2025rethinking] and categorized into outcome models [shao2024deepseekmath] supervising final results and process models [lightman2024lets, wang2024math, zhang2025lessons] supervising intermediate steps. Current video reward models focus on outcome models. Early video reward models adapted image-based approaches [yuan2024instructvideo] or introduced VLMs for quality assessment [wu2024boosting, liu2025improving, he2024videoscore, he2025videoscore2, mougradeo]. However, these outcome-based reward models operate only on near-complete outputs, lacking direct guidance for the critical early denoising stages where fundamental video attributes are established. However, these outcome-based reward models need almost fully denoising stages and lack direct guidance to early denoising stages. LPO [zhang2025diffusion] pioneered using diffusion models as noise-aware latent reward models for image generation, a promising process reward attempt. However, different from image generation, video generation introduces substantially greater complexity. VideoAlign [liu2025improving] briefly mentioned using VGMs as guidance during inference augmentation, while using VGMs as reward models during training remained unexplored. We comprehensively investigate VGMs as latent reward models, offering frame-continuous processing, artifact sensitivity, and timestep-aware rewards aligned with the generation process, with effective integration into the ReFL and other online reinforcement learning methods.
@@ 3 Preliminaries and Feasibility Analysis
Recently, video generation models [kong2024hunyuanvideo, wan2025, gao2025seedance] operate in rectified flow [lipman2023flow, liu2023flow], which establishes a continuous transport between data and noise distributions. Specifically, given clean data $\mathbf{x}_{0}$ sampled from distribution $q(\mathbf{x}_{0})$ and noise $\mathbf{x}_{1}$ from distribution $p(\mathbf{x}_{1})$, the framework defines a time-dependent interpolation path:
@@
Reward Feedback Learning (ReFL) [xu2023imagereward] optimizes diffusion models by backpropagating reward signals through the denoising process. Given a reward model $r_{\phi}$ and generated output $\mathbf{x}_{0}$, ReFL samples a timestep $t$ from the late denoising stage and computes gradients from the reward score. To prevent reward over-optimization, the objective combines reward-based loss with SFT regularization:
@@
VLM fails on noisy inputs (Fig. 3(a)). VideoAlign [liu2025improving], a representative VLM-based video reward model trained on large-scale preference data, exhibits severe performance degradation across different denoising timesteps. When evaluating latents decoded to RGB space, the logit scores fluctuate dramatically between early and late timesteps, revealing poor generalization to high-noise regions.
@@ 4.3 Process Reward Feedback Learning
VideoAlign [liu2025improving] VLM I2V - - - - - 78.83
VideoPhy [bansal2025videophy] VLM I2V - - - - - 77.04
@@ 5.1 Experimental Setup
We collected about 31,000 portrait videos from online sources and generated text prompts using a video caption model. The first frame and text prompt of each video are fed into the Wan2.1-14B-I2V model for inference. The generated videos are labeled by professional annotators based on motion quality as qualified, partially qualified, and unqualified. The partially qualified videos are filtered out to enhance distinctiveness. Detailed annotation and filtration guidelines are provided in the supplementary materials. This process yielded a dataset comprising 24,000 real videos and corresponding generated videos. For reward model, we utilized only the generated videos, and randomly selected 500 samples as the test set (100 for validation, 400 for testing). For video generation, we employed only real videos as SFT training data, and we randomly selected 100 input conditions as our test dataset from the 500-sample test set of the reward model. We also incorporated the existing open-source benchmark VBench [huang2023vbench] and VBench2 [zheng2025vbench2] to ensure a fair comparison with state-of-the-art methods.
@@
We adopt Wan2.1-I2V-14B [wan2025] as our primary baseline. We employ the AdamW optimizer [loshchilov2017decoupled] with learning rates of 1e-5 for query attention and head of PAVRM, 1e-6 for the feature extraction of PAVRM, and 5e-6 for PRFL. We utilize UniPCMultistepScheduler [zhao2023unipc] with 1,000 training steps and 40 inference steps.
@@
For reward model, we select VideoAlign [liu2025improving] and VideoPhy [bansal2025videophy], a state-of-the-art RGB-based reward model. For video generation, we compare against four post-training methods, two offline methods: SFT [liu2023flow] and reward weighted regression (RWR) [liu2025improving], and one online RL method RGB ReFL [lin2025contentv], which only decodes the first frame and uses PickScore [kirstain2023pick] as reward model. See details in supplementary materials.
@@ A.2.2 Open Source Test Set
We incorporated the existing open-source benchmark VBench [huang2023vbench] and VBench2 [zheng2025vbench2] to ensure a fair comparison with state-of-the-art methods. In this paper, we validate the effectiveness of our method in terms of motion quality, which requires the video generation process to be free of distortions, exhibit smooth motion, and comply with physical laws. For the text-to-video (T2V) task, we selected the subject consistency subset from VBench, a total of 72 prompts. Additionally, we employed the human anatomy subset from VBench2, which includes the human anatomy metric with 120 prompts specifically enhanced for the Wan model. For the image-to-video (I2V) task, we selected I2V Subject subset from VBench-I2V (in VBench++ [huang2024vbench++]), a total of 246 prompts.
@@ A.2.4 Baseline Settings
For reward models, we select VideoAlign-MQ [liu2025improving] and VideoPhy-PC [bansal2025videophy], two state-of-the-art VLM-based reward models that excel particularly in assessing motion quality. Both models are employed in a zero-shot manner. The accuracy (Acc) metric is computed by establishing a threshold on the reward scores: videos with scores at or above the threshold are classified as “good”, while those below are classified as “bad”. The accuracy is then calculated against ground-truth labels, where the reported threshold is selected to maximize accuracy on the test set.
@@
SFT [liu2023flow] is a widely adopted and effective post-training technique that offers high computational efficiency. From a reinforcement learning perspective, it can be viewed as an offline, off-policy algorithm, optimizing the loss function defined in Equation 2.
@@
Reward weighted regression (RWR) [liu2025improving] is a prevalent and effective offline, off-policy RL method that has demonstrated success across traditional RL tasks [peng2019advantage], image generation [furuta2024improving], and video generation [liu2025improving]. RWR directly learns from pre-sampled training data treated as experience samples, where a reward model scores each sample to determine its weight in the training loss. The loss function is given by:
@@
For RGB ReFL, we adopt the implementation from ContentV [lin2025contentv], which performs VAE decoding only on the first frame and employs the image reward model PickScore [kirstain2023pick]. The loss function is as follows:
@@ A.2.5 Evaluation
Following the standardized protocol recommended by Wan2.1, we generate evaluation videos at 720P/480P resolution. During the inference phase of video generation models, we maintain a classifier-free guidance (CFG) weight of 5.5. The sampling process employs the UniPCMultistepScheduler [zhao2023unipc] over 40 iterative steps. The early, middle, and late stages of the denoising process correspond to steps 1-13, 14-26, and 27-40, respectively.
@@
Dynamic Degree. To measure generation dynamism, we adopt VBench’s approach using RAFT [teed2020raft] to estimate inter-frame optical flow. We compute the mean of the top 5% flow magnitudes as a static/dynamic threshold, with the final score representing the proportion of non-static videos generated.
Subject Consistency. We adopt VBench’s DINO-based [caron2021emerging] metric to assess subject identity preservation across frames. The consistency score is:
@@
Human Anatomy. We use the VBench metric that they train three ViT-based [xie2022simmim] anomaly detectors for human torso, hands, and faces. Training data includes $\sim$1K real videos (YOLO-World [cheng2024yolo] extracted patches as positives) and $\sim$1K synthetic videos from CogVideo [hong2022cogvideo, yang2024cogvideox] and HunyuanVideo [kong2024hunyuanvideo], plus HumanRefiner [fang2024humanrefiner] negatives, totaling $\sim$150K annotated frames. The score is the percentage of frames without detected anomalies.
I2V Subject Consistency. We use the VBench++ [huang2024vbench++] metric to evaluate input image-to-video subject correspondence. DINOv1 [caron2021emerging] features are extracted from the input image and video frames. The final score combines weighted similarities between the input image and each frame, plus inter-frame similarities, addressing variations in how models handle input images.
@@ A.3.1 The Influence of Training Loss
To assess the robustness of our proposed method against different optimization objectives, we compare the standard binary cross-entropy (BCE) loss with the pairwise Bradley-Terry (BT) loss [bradley1952rank]. Specifically, we construct preference pairs $(x_{\text{win}},x_{\text{lose}})$ by randomly sampling a positive sample (label 1) as $x_{\text{win}}$ and a negative sample (label 0) as $x_{\text{lose}}$. As shown in Table 7, the average accuracy gap between the two objectives is marginal ($0.2\%$). Interestingly, we observe a trade-off across timesteps: BT loss performs better in high-noise regions ($t>0.6$), while BCE demonstrates superior precision in the structure-forming and detailing stages ($t\leq 0.6$). Given that BCE achieves a slightly higher overall average accuracy and eliminates the computational overhead of pair construction, we adopt BCE as our default training objective.
@@ A.3.3 Cross-Model Generalization
To evaluate the generalization capability of PAVRM, we extend our evaluation beyond the source domain. While our reward model is initialized and trained solely on data generated by Wan2.1, we test its performance on samples from two other state-of-the-art video generation models: HunyuanVideo [kong2024hunyuanvideo] and Veo3. The test sets for these models share the same annotation format but specifically focus on human structural deformities (e.g., limb distortions), a common challenge in video generation.

### v3

@@ 1 Introduction
Recent video generation models [21, 19, 36, 10] have demonstrated remarkable visual fidelity, producing photorealistic content across diverse applications [25, 31, 14]. Nevertheless, aligning these models with complex human preferences, particularly concerning motion quality [24], physical plausibility [1], and prompt following [4], remains a fundamental challenge [45]. Reward feedback learning (ReFL) [42, 32, 6] has emerged as a promising way for preference alignment in image generation. Nevertheless, its direct application to video generation encounters critical computational and optimization barriers that fundamentally limit its applicability.
@@
Typical ReFL approaches rely on outcome-based reward models built upon Vision-Language Models (VLMs) [38, 42, 13, 24, 12]. As shown in Fig. 2, such outcome-based reward models focus on videos in RGB space, which imposes three interconnected limitations due to their inherent post-hoc nature for diffusion-based video generation models (VGMs). Initially, the RGB inputs for VLMs require near-complete denoising, which introduces significant evaluation delay and drastically slows down training iterations. Furthermore, this necessity leads to a GPU memory bottleneck, as backpropagation through the VAE decoder for all video frames frequently causes GPU memory overflow. Ultimately, the issue of insufficient supervision occurs because rewards are only applied at the final steps, failing to directly guide early generation stages when structure and motion are formed. Although some recent methods [39, 11] attempt to distribute gradient updates via gradient stopping or trajectory shortcuts, they still depend on fully denoised frames and costly VAE decoding for reward model construction. ContentV [22] bypasses VAE decoding by optimizing only the first frame, yet sacrifices holistic video quality assessment, while DOLLAR [7] employs a VLM-based latent reward model but lacks fine-grained timestep-wise supervision.
Alternatively, VGMs themselves have shown potential as reward sources, offering process-level supervision throughout the denoising trajectory. For instance, LPO [46] pioneered using diffusion models as noise-aware latent reward models for image generation, while VideoAlign [24] briefly mentioned using video generation models for inference-time guidance. However, the use of video generation models as train-time reward models remains largely unexplored. We posit that VGMs possess unique properties for reward modeling: (1) inherent noise-aware feature extraction at arbitrary denoising step, (2) sensitivity to generation artifacts, and (3) native support for full-sequence processing without frame sampling. Through experimental analysis, we verify that the latent features of VGMs are suitable for determining the quality of videos. Yet a critical question remains: how to repurpose VGMs into effective reward models that compress comprehensive spatiotemporal features while maintaining noise-level sensitivity.
@@ 2.1 Visual Reward Feedback Learning
Reward feedback learning has proven effective for improving visual generation quality. Early works [42, 6, 32] finetune diffusion models using differentiable reward signals from human preferences, but face a depth-efficiency dilemma: step-by-step backpropagation through extensive denoising steps incurs prohibitive memory costs, while training only final steps fails to optimize low-level objectives like symmetry. Recent approaches [39, 11] mitigate this through gradient stopping or trajectory-preserving shortcuts. However, these image-focused methods encounter severe computational barriers in video generation due to multi-frame synthesis and gradient-enabled VAE decoding overhead. ContentV [22] optimizes only the first frame but cannot capture comprehensive video quality. Critically, almost existing approaches require pixel-space rewards, necessitating VAE decoding. While DOLLAR [7] leverages a VLM as a latent reward model for distillation tasks to avoid costly VAE decoding, it lacks timestep-wise optimization capability. In this paper, we propose a timestep-wise latent reward model operating directly in latent space, enabling gradient backpropagation at arbitrary timesteps while maintaining computational tractability.
@@ 2.2 Reward Models for Video Generation
Reward models for RLHF [29] are typically trained on human preferences [34] and categorized into outcome models [33] supervising final results and process models [20, 37, 47] supervising intermediate steps. Current video reward models focus on outcome models. Early video reward models adapted image-based approaches [44] or introduced VLMs for quality assessment [40, 24, 13, 12, 28]. However, these outcome-based reward models operate only on near-complete outputs, lacking direct guidance for the critical early denoising stages where fundamental video attributes are established. However, these outcome-based reward models need almost fully denoising stages and lack direct guidance to early denoising stages. LPO [46] pioneered using diffusion models as noise-aware latent reward models for image generation, a promising process reward attempt. However, different from image generation, video generation introduces substantially greater complexity. VideoAlign [24] briefly mentioned using VGMs as guidance during inference augmentation, while using VGMs as reward models during training remained unexplored. We comprehensively investigate VGMs as latent reward models, offering frame-continuous processing, artifact sensitivity, and timestep-aware rewards aligned with the generation process, with effective integration into the ReFL and other online reinforcement learning methods.
@@ 3 Preliminaries and Feasibility Analysis
Recently, video generation models [19, 36, 10] operate in rectified flow [23, 26], which establishes a continuous transport between data and noise distributions. Specifically, given clean data $\mathbf{x}_{0}$ sampled from distribution $q(\mathbf{x}_{0})$ and noise $\mathbf{x}_{1}$ from distribution $p(\mathbf{x}_{1})$, the framework defines a time-dependent interpolation path:
@@
Reward Feedback Learning (ReFL) [42] optimizes diffusion models by backpropagating reward signals through the denoising process. Given a reward model $r_{\phi}$ and generated output $\mathbf{x}_{0}$, ReFL samples a timestep $t$ from the late denoising stage and computes gradients from the reward score. To prevent reward over-optimization, the objective combines reward-based loss with SFT regularization:
@@
VLM fails on noisy inputs (Fig. 3(a)). VideoAlign [24], a representative VLM-based video reward model trained on large-scale preference data, exhibits severe performance degradation across different denoising timesteps. When evaluating latents decoded to RGB space, the logit scores fluctuate dramatically between early and late timesteps, revealing poor generalization to high-noise regions.
@@ 4.3 Process Reward Feedback Learning
VideoAlign [24] VLM I2V - - - - - 78.83
VideoPhy [1] VLM I2V - - - - - 77.04
@@ 5.1 Experimental Setup
We collected about 31,000 portrait videos from online sources and generated text prompts using a video caption model. The first frame and text prompt of each video are fed into the Wan2.1-14B-I2V model for inference. The generated videos are labeled by professional annotators based on motion quality as qualified, partially qualified, and unqualified. The partially qualified videos are filtered out to enhance distinctiveness. Detailed annotation and filtration guidelines are provided in the supplementary materials. This process yielded a dataset comprising 24,000 real videos and corresponding generated videos. For reward model, we utilized only the generated videos, and randomly selected 500 samples as the test set (100 for validation, 400 for testing). For video generation, we employed only real videos as SFT training data, and we randomly selected 100 input conditions as our test dataset from the 500-sample test set of the reward model. We also incorporated the existing open-source benchmark VBench [16] and VBench2 [49] to ensure a fair comparison with state-of-the-art methods.
@@
We adopt Wan2.1-I2V-14B [36] as our primary baseline. We employ the AdamW optimizer [27] with learning rates of 1e-5 for query attention and head of PAVRM, 1e-6 for the feature extraction of PAVRM, and 5e-6 for PRFL. We utilize UniPCMultistepScheduler [48] with 1,000 training steps and 40 inference steps.
@@
For reward model, we select VideoAlign [24] and VideoPhy [1], a state-of-the-art RGB-based reward model. For video generation, we compare against four post-training methods, two offline methods: SFT [26] and reward weighted regression (RWR) [24], and one online RL method RGB ReFL [22], which only decodes the first frame and uses PickScore [18] as reward model. See details in supplementary materials.
@@ A.2.2 Open Source Test Set
We incorporated the existing open-source benchmark VBench [16] and VBench2 [49] to ensure a fair comparison with state-of-the-art methods. In this paper, we validate the effectiveness of our method in terms of motion quality, which requires the video generation process to be free of distortions, exhibit smooth motion, and comply with physical laws. For the text-to-video (T2V) task, we selected the subject consistency subset from VBench, a total of 72 prompts. Additionally, we employed the human anatomy subset from VBench2, which includes the human anatomy metric with 120 prompts specifically enhanced for the Wan model. For the image-to-video (I2V) task, we selected I2V Subject subset from VBench-I2V (in VBench++ [17]), a total of 246 prompts.
@@ A.2.4 Baseline Settings
For reward models, we select VideoAlign-MQ [24] and VideoPhy-PC [1], two state-of-the-art VLM-based reward models that excel particularly in assessing motion quality. Both models are employed in a zero-shot manner. The accuracy (Acc) metric is computed by establishing a threshold on the reward scores: videos with scores at or above the threshold are classified as “good”, while those below are classified as “bad”. The accuracy is then calculated against ground-truth labels, where the reported threshold is selected to maximize accuracy on the test set.
@@
SFT [26] is a widely adopted and effective post-training technique that offers high computational efficiency. From a reinforcement learning perspective, it can be viewed as an offline, off-policy algorithm, optimizing the loss function defined in Equation 2.
@@
Reward weighted regression (RWR) [24] is a prevalent and effective offline, off-policy RL method that has demonstrated success across traditional RL tasks [30], image generation [9], and video generation [24]. RWR directly learns from pre-sampled training data treated as experience samples, where a reward model scores each sample to determine its weight in the training loss. The loss function is given by:
@@
For RGB ReFL, we adopt the implementation from ContentV [22], which performs VAE decoding only on the first frame and employs the image reward model PickScore [18]. The loss function is as follows:
@@ A.2.5 Evaluation
Following the standardized protocol recommended by Wan2.1, we generate evaluation videos at 720P/480P resolution. During the inference phase of video generation models, we maintain a classifier-free guidance (CFG) weight of 5.5. The sampling process employs the UniPCMultistepScheduler [48] over 40 iterative steps. The early, middle, and late stages of the denoising process correspond to steps 1-13, 14-26, and 27-40, respectively.
@@
Dynamic Degree. To measure generation dynamism, we adopt VBench’s approach using RAFT [35] to estimate inter-frame optical flow. We compute the mean of the top 5% flow magnitudes as a static/dynamic threshold, with the final score representing the proportion of non-static videos generated.
Subject Consistency. We adopt VBench’s DINO-based [3] metric to assess subject identity preservation across frames. The consistency score is:
@@
Human Anatomy. We use the VBench metric that they train three ViT-based [41] anomaly detectors for human torso, hands, and faces. Training data includes $\sim$1K real videos (YOLO-World [5] extracted patches as positives) and $\sim$1K synthetic videos from CogVideo [15, 43] and HunyuanVideo [19], plus HumanRefiner [8] negatives, totaling $\sim$150K annotated frames. The score is the percentage of frames without detected anomalies.
I2V Subject Consistency. We use the VBench++ [17] metric to evaluate input image-to-video subject correspondence. DINOv1 [3] features are extracted from the input image and video frames. The final score combines weighted similarities between the input image and each frame, plus inter-frame similarities, addressing variations in how models handle input images.
@@ A.3.1 The Influence of Training Loss
To assess the robustness of our proposed method against different optimization objectives, we compare the standard binary cross-entropy (BCE) loss with the pairwise Bradley-Terry (BT) loss [2]. Specifically, we construct preference pairs $(x_{\text{win}},x_{\text{lose}})$ by randomly sampling a positive sample (label 1) as $x_{\text{win}}$ and a negative sample (label 0) as $x_{\text{lose}}$. As shown in Table 7, the average accuracy gap between the two objectives is marginal ($0.2\%$). Interestingly, we observe a trade-off across timesteps: BT loss performs better in high-noise regions ($t>0.6$), while BCE demonstrates superior precision in the structure-forming and detailing stages ($t\leq 0.6$). Given that BCE achieves a slightly higher overall average accuracy and eliminates the computational overhead of pair construction, we adopt BCE as our default training objective.
@@ A.3.3 Cross-Model Generalization
To evaluate the generalization capability of PAVRM, we extend our evaluation beyond the source domain. While our reward model is initialized and trained solely on data generated by Wan2.1, we test its performance on samples from two other state-of-the-art video generation models: HunyuanVideo [19] and Veo3. The test sets for these models share the same annotation format but specifically focus on human structural deformities (e.g., limb distortions), a common challenge in video generation.